# Polarized and unpolarized gluon PDFs: generative machine learning applications for lattice QCD matrix elements at short distance and large momentum

### v1

@@ IV Application of ML algorithms to the lattice data
The random forest (RF) regressor is the simplest ensemble method, in which each regressor is a decision tree trained on a random sample $S_{i}$ of the data, and each tree has a unit weight in the ensemble [61]. Therefore, random forests are computationally efficient and have low variance. The number of trees and the depth of the tree are two hyperparameters of an RF regressor. By selecting a random subset of features for each tree and training each tree on a bootstrapped dataset sample, this randomness helps prevent overfitting and decorrelates the individual trees, resulting in a more accurate and stable predictor.
As proposed in [62], the gradient boosted decision tree (GBT) [63] is another tree-based ensemble method where several decision trees are used as weak learners. Together, all the weak learners form a strong and effective regressor. Initially, each sample is weighted equally, and then the first weak regressor trains the first sample, and so on. After learning, one reduces the weight of the correctly predicted samples and increases the weight of the mistaken samples. Residuals are computed from the mistaken samples, and a weak regressor is trained based on the previous weak regressor’s residual error. In this way, GBT can reach the prediction target by decreasing the residual error in the training process.
Extreme gradient boost (XGB) regressor [64] is an improved version of the GBT regressor regarding computing speed, generalization, and scalability. Its efficient implementation uses parallel and distributed computing, allowing faster training times than traditional gradient-boosting algorithms. One of XGB’s critical features is its ability to handle missing data and outliers effectively. XGB enhances model generalization, mitigates overfitting, and provides robustness in noisy or incomplete datasets by employing a sparsity-aware algorithm and incorporating regularization terms.
@@ IV.2 Experimental results
LQCD dataset associated with the polarized gluon distribution consists of a vector of $1901$ data points ($\Delta\mathfrak{M}$) for each pair of momentum $p_{z}$ and $\omega$. 1901 is the number of configurations used in the LQCD calculations [36, 37]. By each pair of momentum, we refer to the relation between two-$p_{z}$ datasets in Eq. (7). We computed an aggregate dataset $\mathcal{D}_{r}$ such that $\mathcal{D}_{r}=\{(p,w,\mu_{r}(w),\sigma_{r}(w)):\mu_{r}(w)\leftarrow{\rm mean}(\Delta\mathfrak{M}(w)),\sigma_{r}(w)\leftarrow{\rm std.deviation}(\Delta\mathfrak{M}(w)),p\in p_{z},w\in\omega\}$. Fig. 2 shows the $\Delta\mathfrak{M}(\omega)$ vs $\omega$, and the points are simply connected for each $p_{z}$. $\Delta\mathfrak{M}(\omega)=0$ when $\omega=0$. When $\omega$ increases, different curves (for different $p_{z}$) have different contributions from the contamination term, as seen from Eq. (4). $\Delta\mathfrak{M}$ curves move towards the phenomenological ITD [43] as $p_{z}$ increases and show features of decreasing around $\omega\geq 6$ at $p_{z}=2.46$ GeV within uncertainty. The challenge of removing the contamination term $(m_{p}^{2}/p_{z}^{2})\omega\Delta{\mathcal{M}}_{pp}$ and obtaining a nonzero signal can be understood from the smallness in the phenomenological gluon helicity ITD, as noted in [43, 37]. The pattern of the curve outside the available lattice data is unclear. Due to the small number of observations, any ML model will struggle with underfitting and the uncertainty of the extrapolated data will be large.
@@
To overcome this limitation, we incorporated physics constraints and knowledge from the phenomenological ITD from [43] (or alternatively can be used from [53]) along with the aggregate dataset ($\mathcal{D}_{r}$) to generate synthetic data. We have used two phenomenological ITDs from different fits in [43], as shown in Fig. 3. The magnitude of the ML-generated data is not affected by those of the phenomenological ITDs. The knowledge of the phenomenological ITDs outside the LQCD data domain helps to determine the shape of the ML-generated ITD at large $\omega$ after eliminating the contamination term. Also, as discussed in [37], within the current statistics, the matching effects on $\Delta\mathfrak{M}_{g}(\omega)$ in the $\overline{\rm{MS}}$-scheme is marginal within uncertainty. One can convert the phenomenological ITD data to the reduced pseudo-ITD data using the matching formula in [39] and verify this. Therefore, we assume the shape of the phenomenological ITD is similar to that of the $\Delta\mathfrak{M}_{g}(\omega)$ where LQCD data is unavailable. One can numerically check that synthetic data using either the fit-1 or fit-2 curves in Fig. 3 results in the same curves from the lattice data. For example, as we will see, the two phenomenological curves have maximum values, $\Delta\mathcal{I}_{g}\approx 0.4$ and $0.2$. In contrast, the contamination-free $\Delta\mathfrak{M}_{g}$ will have a maximum value of around $0.12$ by fitting the LQCD data along with the synthetic data.
@@
Figure 3: Phenomenological shapes of $\Delta\mathcal{I}_{g}(\omega,\mu)$ with respect to $\omega$ from [43] in the $\overline{\rm{MS}}$ scheme at $\mu=2$ GeV. As discussed in the text, we focus here on the central lines of two different fit results from [43] to determine the shape of the ITD.
In the following, we will present a specific case of ML, the gradient descent algorithm [65]. Fig. 3 shows the pattern of the phenomenological $\Delta\mathcal{I}_{g}(\omega)$, which is used to derive $\Delta\mathfrak{M}(\omega)$ using the gradient descent algorithm. Gradient descent is an optimization algorithm that finds the minimum or maximum in the loss function. A loss function must be either convex or concave to have a minimum or maximum value, respectively. In our solution, we have $\Delta\mathfrak{M}_{g}(\omega)$ from Eq. (IV), and $\Delta\mathcal{I}_{g}(\omega)$ from Fig. 3. The convex loss function is defined as $\mathcal{L}\equiv(\Delta\mathfrak{M}_{g}(\omega)-\Delta\mathcal{I}_{g}(\omega))^{2}$. Derivation of the loss function concerning $\Delta\mathfrak{M}_{g}(\omega)$ denoted by $\partialderivative{\mathcal{L}}{\Delta\mathfrak{M}_{g}(\omega)}$ indicates a slope at $\Delta\mathfrak{M}_{g}(\omega)$ on the loss curve. We update the $\Delta\mathfrak{M}_{g}(\omega)$ with respect to the derivative $\Delta\mathfrak{M}_{g}(\omega)=\Delta\mathfrak{M}_{g}(\omega)-\eta\partialderivative{\mathcal{L}}{\Delta\mathfrak{M}_{g}(\omega)}$. This ensures the shift of $\Delta\mathfrak{M}_{g}(\omega)$ towards the minimum point. The pace or step size of the update depends on $\eta$, which is called the learning rate. Algorithm 1 shows the details to estimate $\Delta\mathfrak{M}_{g}(\omega)$ by generating the synthetic data $\mathcal{D}_{g}$.
@@ 22      $\mathcal{D}_{g}\leftarrow Append(\mathcal{D}_{g},\mathcal{G})$
    We adjusted the initialization of $\Delta\mathfrak{M}_{i}$ and $\Delta\mathfrak{M}_{j}$ to achieve the best overlapping between the real and the estimated values (shown in Fig. 5). To optimize the overlap and adhere to the relation between two $p_{z}$ datasets, it becomes apparent that only $z\leq 4a(=0.36~{}{\rm fm})$ LQCD matrix elements can be incorporated into the fit while satisfying Eq. (IV).  [^1] This observation aligns with the findings in [37], indicating that LQCD matrix elements at large physical separations may involve substantial power corrections, rendering them unsuitable for determining the light-cone correlation function associated with the gluon helicity PDF. Further discussion on this topic follows in the subsequent section of the manuscript.
@@
$\mathbb{P}(\Delta\mathfrak{M}_{\ell}|\omega_{\ell},\Delta\mathfrak{M}_{\ell-1},\omega_{\ell-1},\dots,\Delta\mathfrak{M}_{\ell-k},\omega_{\ell-k}),$ (14)
@@
We used the curve fitting method to find the best-fitted model for mean and standard deviation of $\mathfrak{M}(\omega)$ and the fitted model is used to generate mean and standard deviation for new $\omega$. We used dose-response [67] based curve fitting models to fit mean values. 10 models are used to fit a curve, and the best-fitted model is selected using AIC (Akaike Information Criterion) log-likelihood. Fig. 10 shows the best-fitted curve for each $p_{z}$.
@@ V Discussion on the ML-generated data and results
Machine learning is a data-dependent algorithm where the decision process depends on the data pattern. Sufficient data is a necessary condition to reveal and understand the data pattern. Because there is insufficient data to reveal the pattern, we generated synthetic data to mitigate the gap and achieve the experimental objective. To derive synthetic data, we deployed ML models to extract a pattern for the mean and standard deviation of $\Delta\mathfrak{M}(\omega)$ for each $\omega$ derived from LQCD data. The well-fitted ML model extrapolates mean and standard deviation for unknown and large values of $\omega$; therefore, for each $\omega$, we have a collection of $(\mu_{g},\sigma_{g})$. $N$ synthetic data are picked from a normal distribution, $\mathcal{N}(\mu_{g},\sigma_{g})$. For each $\omega$, there are $N$ values of $\Delta\mathfrak{M}(\omega)$ used to compute the uncertainty.
The ML model on the standard deviation of the real data for each $\omega$ controls the magnitude of the uncertainty. For the polarized data, a linear model fits the polarized data (in Fig. 6). The slope of the line indicates the rate of the change of the magnitude of the standard deviation for the change of the $\omega$; a steeper slope means the value of the standard deviation will increase a lot for a small increase of the value of $\omega$. Alternatively, a gradual slope means a small standard deviation increase for a large increase of the $\omega$. The slope value of Fig. 6 indicates that $p_{z}=1$ has a higher slope than $p_{z}=6$. Therefore, we noticed smaller uncertainty for $p_{z}=6$ compared to $p_{z}=1$ ML-generated dataset.

### v2

@@ IV Application of ML algorithms to the lattice data
As we will demonstrate in the following, due to the small number of observations and limited data in the $\omega\in[0,7]$ region, the pattern of the curves outside the available lattice data, along with the constraint in Eq. (7) across different $p_{z}$ datasets, suggests that the ML model fits can benefit from the phenomenological polarized gluon ITD at large $\omega$. In particular, the term “physics-informed ML” refers to the combination of the constraint in Eq. (7), which we enforce across different $p_{z}$ datasets with minimal power corrections, in addition to the contamination term $(m_{p}^{2}/p_{z}^{2})\omega\Delta\mathcal{M}_{pp}$ and the phenomenological ITD at large $\omega$. We note that no such constraints will be applied when fitting the unpolarized gluon ITD data. Therefore, by physics-informed ML, we exclusively refer to the case of the polarized gluon distribution.
@@
Machine learning is a data-driven algorithm that uses different statistical processes to extract implicit patterns from high dimensional high volume data. The pattern is used for inference or association with unseen data. Based on the data demography, ML algorithms are divided into three categories: a) supervised learning, b) unsupervised learning, and c) reinforcement learning. Supervised learning algorithms are used for the data where each observation is labeled. The unsupervised learning algorithms are used for the data where the label for each observation is missing. Reinforcement learning gathers data from its environment, processes it, and generates the next action. For instance, a drone collects images from its surroundings, processes them, and finds a path to move forward, backward, and upward or downward. In labeled data, the label can be categorical or real-valued. The ML algorithm used for categorical labeled data is called the classification algorithm, and the algorithm used for real-valued labeled data is called the regression algorithm.
Figure 2: The figure shows the necessary components for the machine learning process. The process starts with data. The feature engineering process extracts important features and prepares data for the next level of operations. The data is split into two disjoint subsets: training and testing data. The training data ($x_{tr}$) is used to train a machine learning model, and the testing data ($x_{ts}$) is used to evaluate the performance of the trained model.
Fig. 2 shows the general pipeline of any ML algorithm. Data is the most essential element of any ML process. Most of the time, the collected data needs pre-processing, which is called data-wrangling. In the data-wrangling process, one might remove missing and inconsistent data. This step also includes feature engineering, which aims to consider only the important features and data transformation. One good practice is normalizing the features’ values to make them comparable. The preprocessed data is split into two disjoint subsets: training and testing data. Based on the volume of data, training data is further subdivided into two disjoint subsets named training and validation. The split proportion is user-defined, but the general convention is to $80\%$ training and $20\%$ testing (or $10\%$ testing and $10\%$ validation).
The training data is used to train an ML model. A model is a function of input variables $X$ and unknown variables $\theta$. It is referred to as parameter because the model learns the values of $\theta$ from data. Another set of variables is required to manipulate the ML system called hyperparameter, for instance, learning rate. These kinds of variables are like settings of an ML algorithm while training. Once the model is trained, we test data with the same parameter and hyperparameter configuration.
For instance, a decision tree (DT) is a rule-based machine learning algorithm where it divides the data in a tree fashion but visually represents it in a reverse way. It uses Gini index Daniya et al. (2020) to choose the best feature to split the data. The root node is a feature that has the highest Gini index, and each of the left nodes holds a decision. As the tree grows, it holds all the information of the training data and may fail to classify some instances for testing data. To avoid this overfitting scenario, a hyperparameter named “maxheight” controls the tree’s growth. Like all hyperparameters, the user can control this variable and prune the tree to control overfitting.
Figure 3: The figure shows the necessary components for an ensemble machine learning process. The process starts with data. The feature engineering process extracts important features and prepares data for the next level of operations. The data is split into two disjoint subsets: training and testing data. The training dataset is split into $k$ sample datasets. Based on the ensemble model algorithm, each subsample of data trains a weak learner and finally aggregates the outputs (decisions) to make a combined decision.
In many scenarios, combined decisions from multiple models have better prediction probability than a single model. FIG. 3 shows the basic architecture of an ensemble model. Each model in the ensemble model is a weak learner, and the combined decision of all weak learners produces a strong decision. A weak learner is a machine learning model with better prediction power than random choice. Ensemble models are categorized into three parts: a) Parallel, b) Sequential, and c) Composite. Bagging and Voting ensemble models are parallel models where many weak learners are trained parallelly on different data samples. The combined decision is used as the model’s decision. Random forest is a popular bagging algorithm. Boosting algorithms like “Gradient boosting” or “XGBoost” also use many weak learners, but they train the model sequentially.
@@
The random forest (RF) regressor is the simplest ensemble method, in which each regressor is a decision tree trained on a random sample $S_{i}$ of the data, and each tree has a unit weight in the ensemble Breiman (1). Therefore, random forests are computationally efficient and have low variance. The number of trees and the depth of the tree are two hyperparameters of an RF regressor. By selecting a random subset of features for each tree and training each tree on a bootstrapped dataset sample, this randomness helps prevent overfitting and decorrelates the individual trees, resulting in a more accurate and stable predictor.
As proposed in Friedman (2001), the gradient boosted decision tree (GBT) Boehmke and Greenwell (2019) is another tree-based ensemble method where several decision trees are used as weak learners. Together, all the weak learners form a strong and effective regressor. Initially, each sample is weighted equally, and then the first weak regressor trains the first sample, and so on. After learning, one reduces the weight of the correctly predicted samples and increases the weight of the mistaken samples. Residuals are computed from the mistaken samples, and a weak regressor is trained based on the previous weak regressor’s residual error. In this way, GBT can reach the prediction target by decreasing the residual error in the training process.
Extreme gradient boost (XGB) regressor Chen and Guestrin (2016) is an improved version of the GBT regressor regarding computing speed, generalization, and scalability. Its efficient implementation uses parallel and distributed computing, allowing faster training times than traditional gradient-boosting algorithms. One of XGB’s critical features is its ability to handle missing data and outliers effectively. XGB enhances model generalization, mitigates overfitting, and provides robustness in noisy or incomplete datasets by employing a sparsity-aware algorithm and incorporating regularization terms.
@@ IV.2 Experimental results
LQCD dataset associated with the polarized gluon distribution consists of a vector of $1901$ data points ($\Delta\mathfrak{M}$) for each pair of momentum $p_{z}$ and $\omega$. 1901 is the number of configurations used in the LQCD calculations Egerer et al. (2022); Khan et al. (2023). By each pair of momentum, we refer to the relation between two-$p_{z}$ datasets in Eq. (7). We computed an aggregate dataset $\mathcal{D}_{r}$ such that $\mathcal{D}_{r}=\{(p,w,\mu_{r}(w),\sigma_{r}(w)):\mu_{r}(w)\leftarrow{\rm mean}(\Delta\mathfrak{M}(w)),\sigma_{r}(w)\leftarrow{\rm std.deviation}(\Delta\mathfrak{M}(w)),p\in p_{z},w\in\omega\}$. Fig. 4 shows the $\Delta\mathfrak{M}(\omega)$ vs $\omega$, and the points are simply connected for each $p_{z}$. $\Delta\mathfrak{M}(\omega)=0$ when $\omega=0$. When $\omega$ increases, different curves (for different $p_{z}$) have different contributions from the contamination term, as seen from Eq. (4). $\Delta\mathfrak{M}$ curves move towards the phenomenological ITD Sufian et al. (2021) as $p_{z}$ increases and show features of decreasing around $\omega\geq 6$ at $p_{z}=2.46$ GeV within uncertainty. The challenge of removing the contamination term $(m_{p}^{2}/p_{z}^{2})\omega\Delta{\mathcal{M}}_{pp}$ and obtaining a nonzero signal can be understood from the smallness in the phenomenological gluon helicity ITD, as noted in Sufian et al. (2021); Khan et al. (2023). The pattern of the curve outside the available lattice data is unclear. Due to the small number of observations, any ML model will struggle with underfitting and the uncertainty of the extrapolated data will be large.
@@
To overcome this limitation, we incorporated physics constraints and knowledge from the phenomenological ITD from Sufian et al. (2021) (or alternatively can be used from Ball, Richard D. et al. (2017)) along with the aggregate dataset ($\mathcal{D}_{r}$) to generate synthetic data. We have used two phenomenological ITDs from different fits in Sufian et al. (2021), as shown in Fig. 5. The magnitude of the ML-generated data is not affected by those of the phenomenological ITDs. The knowledge of the phenomenological ITDs outside the LQCD data domain helps to determine the shape of the ML-generated ITD at large $\omega$ after eliminating the contamination term. Also, as discussed in Khan et al. (2023), within the current statistics, the matching effects on $\Delta\mathfrak{M}_{g}(\omega)$ in the $\overline{\rm{MS}}$-scheme is marginal within uncertainty. One can convert the phenomenological ITD data to the reduced pseudo-ITD data using the matching formula in Balitsky et al. (2022) and verify this. Therefore, we assume the shape of the phenomenological ITD is similar to that of the $\Delta\mathfrak{M}_{g}(\omega)$ where LQCD data is unavailable. One can numerically check that synthetic data using either the fit-1 or fit-2 curves in Fig. 5 results in the same curves from the lattice data. For example, as we will see, the two phenomenological curves have maximum values, $\Delta\mathcal{I}_{g}\approx 0.4$ and $0.2$. In contrast, the contamination-free $\Delta\mathfrak{M}_{g}$ will have a maximum value of around $0.12$ by fitting the LQCD data along with the synthetic data.
@@
Figure 5: Phenomenological shapes of $\Delta\mathcal{I}_{g}(\omega,\mu)$ with respect to $\omega$ from Sufian et al. (2021) in the $\overline{\rm{MS}}$ scheme at $\mu=2$ GeV. As discussed in the text, we focus here on the central lines of two different fit results from Sufian et al. (2021) to determine the shape of the ITD.
In the following, we will present a specific case of ML, the gradient descent algorithm Ruder (2016). Fig. 5 shows the pattern of the phenomenological $\Delta\mathcal{I}_{g}(\omega)$, which is used to derive $\Delta\mathfrak{M}(\omega)$ using the gradient descent algorithm. Gradient descent is an optimization algorithm that finds the minimum or maximum in the loss function. A loss function must be either convex or concave to have a minimum or maximum value, respectively. In our solution, we have $\Delta\mathfrak{M}_{g}(\omega)$ from Eq. (IV), and $\Delta\mathcal{I}_{g}(\omega)$ from Fig. 5. The convex loss function is defined as $\mathcal{L}\equiv(\Delta\mathfrak{M}_{g}(\omega)-\Delta\mathcal{I}_{g}(\omega))^{2}$. Derivation of the loss function concerning $\Delta\mathfrak{M}_{g}(\omega)$ denoted by $\partialderivative{\mathcal{L}}{\Delta\mathfrak{M}_{g}(\omega)}$ indicates a slope at $\Delta\mathfrak{M}_{g}(\omega)$ on the loss curve. We update the $\Delta\mathfrak{M}_{g}(\omega)$ with respect to the derivative $\Delta\mathfrak{M}_{g}(\omega)=\Delta\mathfrak{M}_{g}(\omega)-\eta\partialderivative{\mathcal{L}}{\Delta\mathfrak{M}_{g}(\omega)}$. This ensures the shift of $\Delta\mathfrak{M}_{g}(\omega)$ towards the minimum point. The pace or step size of the update depends on $\eta$, which is called the learning rate. Algorithm 1 shows the details to estimate $\Delta\mathfrak{M}_{g}(\omega)$ by generating the synthetic data $\mathcal{D}_{g}$.
@@ 22      $\mathcal{D}_{g}\leftarrow Append(\mathcal{D}_{g},\mathcal{G})$
    We adjusted the initialization of $\Delta\mathfrak{M}_{i}$ and $\Delta\mathfrak{M}_{j}$ to achieve the best overlapping between the real and the estimated values (shown in Fig. 7). To optimize the overlap and adhere to the relation between two $p_{z}$ datasets, it becomes apparent that only $z\leq 4a(=0.36~{}{\rm fm})$ LQCD matrix elements can be incorporated into the fit while satisfying Eq. (IV).  [^1] This observation aligns with the findings in Khan et al. (2023), indicating that LQCD matrix elements at large physical separations may involve substantial power corrections, rendering them unsuitable for determining the light-cone correlation function associated with the gluon helicity PDF. Further discussion on this topic follows in the subsequent section of the manuscript.
@@
$\mathbb{P}_{r}(\Delta\mathfrak{M}_{\ell}|\omega_{\ell},\Delta\mathfrak{M}_{\ell-1},\omega_{\ell-1},\dots,\Delta\mathfrak{M}_{\ell-k},\omega_{\ell-k}),$ (14)
@@
We used the curve fitting method to find the best-fitted model for mean and standard deviation of $\mathfrak{M}(\omega)$ and the fitted model is used to generate mean and standard deviation for new $\omega$. We used dose-response Ritz et al. (2015) based curve fitting models to fit mean values. 10 models are used to fit a curve, and the best-fitted model is selected using AIC (Akaike Information Criterion) log-likelihood. Fig. 12 shows the best-fitted curve for each $p_{z}$.
@@ V Discussion on the ML-generated data and results
Machine learning is a data-dependent algorithm where the decision process depends on the quality and quantity of the data. Sufficient data is necessary for an ML model to learn the hidden pattern in the data. Due to the lack of these properties in our experimental data, we generated synthetic data to mitigate the gap and achieve the experimental objective. The synthetic data $X$ is randomly generated from a normal distribution $X\sim\mathcal{N}(\mu,\sigma)$ where mean $\mu$ and standard deviation $\sigma$ are predicted using two different machine learning models. The mean is predicted from an ensemble model, and the standard deviation is predicted from a linear regression model. There is uncertainty in each machine learning model to compute mean and standard deviation, which imposes another layer of uncertainty when deriving synthetic data from a normal distribution constructed using that mean and standard deviation. We observed high variations of the error bar for the XGB model in FIG. 14 compared to the RF and GBT models. This is because of the XGB model’s limitation in handling noisy synthetic datasets. This limitation of XGB is reflected in Fig. 14. Further investigations with more precise data are needed to quantitatively assess the systematic uncertainties associated with these ML models.
The choice of a proper ML model is critical and mainly depends on the data. The volume and dimension of data exclude the choice of using any deep learning model and leave us to use classical ML models. We tried the linear regression model, but the model did not converge on our data. The support vector regressor with a polynomial kernel also suffers the same issue, but the radial basis function (RBF) kernel cannot outperform ensemble models. Therefore, we stick to ensemble models for our experiment.
The ML model on the standard deviation of the real data for each $\omega$ controls the magnitude of the uncertainty. For the polarized data, a linear model fits the polarized data (in Fig. 8). The slope of the line indicates the rate of the change of the magnitude of the standard deviation for the change of the $\omega$; a steeper slope means the value of the standard deviation will increase a lot for a small increase of the value of $\omega$. Alternatively, a gradual slope means a small standard deviation increase for a large increase of the $\omega$. The slope value of Fig. 8 indicates that $p_{z}=1$ has a higher slope than $p_{z}=6$. Therefore, we noticed smaller uncertainty for $p_{z}=6$ than $p_{z}=1$ ML-generated dataset.